# Beyond Synthetic Benchmarks: Evaluating LLM Performance on Real-World Class-Level Code Generation

### v1

@@ Abstract.
Large language models (LLMs) have advanced code generation at the function level, yet their ability to produce correct class-level implementations in authentic software projects remains poorly understood. This work introduces a novel benchmark derived from open-source repositories, comprising real-world classes divided into seen and unseen partitions to evaluate generalization under practical conditions. The evaluation examines multiple LLMs under varied input specifications, retrieval-augmented configurations, and documentation completeness levels.
Results reveal a stark performance disparity: LLMs achieve 84–89% correctness on established synthetic benchmarks but only 25–34% on real-world class tasks, with negligible differences between familiar and novel codebases. Comprehensive docstrings yield modest gains of 1–3% in functional accuracy, though statistical significance is rare. Retrieval-augmented generation proves most effective with partial documentation, improving correctness by 4–7% by supplying concrete implementation patterns absent from specifications. Error profiling identifies AttributeError, TypeError, and AssertionError as dominant failure modes (84% of cases), with synthetic tests overemphasizing assertion issues and real-world scenarios highlighting type and attribute mismatches. Retrieval augmentation reduces logical flaws but can introduce dependency conflicts.
The benchmark and analysis expose critical limitations in current LLM capabilities for class-level engineering, offering actionable insights for enhancing context modelling, documentation strategies, and retrieval integration in production code assistance tools.
@@ 1. Introduction
Large language models (LLMs) have demonstrated remarkable progress in code generation, enabling tools like GitHub Copilot and Amazon CodeWhisperer to assist developers in writing functions, classes, and even entire modules (chen2021evaluating, ; tao2025retrieval, ). However, existing benchmarks such as CoderEval (yu2024codereval, ) and ClassEval (du2023classeval, ) primarily focus on either function-level generation or manually crafted class-level tasks. These benchmarks do not fully capture the complexities of real-world software projects, where developers frequently encounter interdependent classes, project-specific patterns, and the need to generalize to unseen codebases. This gap raises a critical question: How well do LLMs generate class-level code for real-world projects, particularly when faced with unseen contexts? Addressing this question is essential for advancing LLM-assisted programming tools and ensuring their reliability in practical software development scenarios.
@@
Current research on LLM-based code generation has largely centred on function-level tasks, with benchmarks like HumanEval (chen2021evaluating, ) and MBPP (austin2021program, ) evaluating standalone functions. While ClassEval (du2023classeval, ) introduced class-level evaluation, its manually crafted tasks do not reflect the intricacies of real-world projects, such as cross-class dependencies and evolving codebases. Recent work on retrieval-augmented generation (RAG) for code (yang2025empirical, ; tao2025retrieval, ) has shown promise in improving LLM performance by incorporating relevant context, but its application to class-level generation in real-world settings remains underexplored. Furthermore, studies on repository-level code generation (tao2025retrieval, ) highlight the need for benchmarks that capture project-level dependencies, which are absent in existing evaluations.
@@
We address the issues above by curating a new benchmark of real-world class-level code from open-source projects, categorized as ‘seen’ and ‘unseen’. Our evaluation framework assesses LLM performance across different docstring variants, generation strategies, and RAG-augmented setups. We systematically analyze the impact of docstrings, the effectiveness of RAG, and the types of errors that arise in generated code, providing insights into the strengths and limitations of current LLMs. More specifically, we investigate the following research questions:
@@
  • RQ4 (Error Analysis): What are the most common errors in LLM-generated class-level code, and how do they differ between seen and unseen contexts? We find that AttributeError, TypeError, and AssertionError account for 84% of all failures. Synthetic benchmarks show assertion-dominated failures while real-world code shows attribute-and-type-dominated failures, and RAG operates through error substitution, reducing logic failures while occasionally introducing dependency errors.
@@ 2.1. Large Language Models for Code Generation
Large Language Models (LLMs) have achieved remarkable success in code generation, enabling tools like GitHub Copilot and Amazon CodeWhisperer to assist developers in writing functions, classes, and even entire modules (chen2021evaluating, ; austin2021program, ). Recent works highlight the rapid evolution of LLMs for code-related tasks, including code completion, translation, and repair (jiang2024survey, ; bouzenia2024repairagent, ). Models such as GPT-4, CodeGen, and WizardCoder have demonstrated strong performance on benchmarks like HumanEval and MBPP, but their effectiveness in real-world scenarios—particularly for class-level code generation—remains underexplored (du2023classeval, ).
@@ 2.2. Benchmarks for Code Generation
Existing benchmarks for code generation primarily focus on function-level tasks. For example, HumanEval (chen2021evaluating, ) and MBPP (austin2021program, ) evaluate standalone functions, while CoderEval (yu2024codereval, ) introduces non-standalone functions but still operates at the function level. ClassEval (du2023classeval, ) is the first benchmark to evaluate class-level code generation, but its tasks are manually crafted and do not reflect the complexities of real-world projects, such as cross-class dependencies and project-specific patterns.
@@ 2.3. Retrieval-Augmented Generation for Code
Retrieval-Augmented Generation (RAG) has emerged as a promising approach to enhance LLM performance by incorporating relevant context during generation (yang2025empirical, ; tao2025retrieval, ). Recent studies have explored RAG for code generation, demonstrating its potential to improve accuracy and contextual relevance (li2025coderag, ). For instance, CodeRAG (li2025coderag, ) integrates retrieval into the reasoning process, allowing LLMs to interact with source code snippets from repositories. However, most RAG studies focus on function-level or file-level tasks, and their application to class-level generation in real-world settings remains limited (tao2025retrieval, ).
@@ 2.4. Docstrings and Prompt Engineering
Docstrings play a crucial role in code generation, as they provide natural language descriptions of code functionality. Studies have shown that the quality and presence of docstrings significantly impact LLM performance (yang2024less, ; ronanki2025prompt, ). For example, ClassEval (du2023classeval, ) includes human-labelled docstrings to mitigate data leakage and improve prompt quality. Recent work on docstring compression (yang2024less, ) further highlights the importance of concise and informative docstrings for effective code generation. Prompt engineering techniques, such as chain-of-thought and few-shot prompting (yuenprompting, ), have also been explored to enhance LLM performance in code-related tasks (tony2025prompting, ; sahoo2024systematic, ).
@@ 3.1. Dataset Curation
To evaluate the performance of LLMs on real-world class-level code generation, we introduce RealClassEval, a dataset comprising two versions of Python classes: pre-cutoff and post-cutoff. Each version contains 200 classes mined from real-world GitHub open-source projects. The pre-cutoff data is sourced exclusively from the CodeSearchNet dataset (husain2019codesearchnet, ). Even though we cannot guarantee that this dataset was indeed part of the training data of the LLMs we investigated in this research, we believe that it is a safe assumption to make, given the popularity of this dataset. Furthermore, several existing LLMs (wang2023codet5+, ; li2023starcoder, ) that are not part of this study have explicitly mentioned the use of CodeSearchNer in their training corpus. The post-cutoff data consists of repositories created after March 31, 2025  [^1] All LLMs under investigation were either released before this date or their knowledge cut-off dates are before this date., guaranteeing these classes are unseen by the evaluated LLMs.
We retain only engineered projects (munaiah2017curating, ), adhering to the filtering criteria outlined in (xiao2025self, ). Using Understand™ (scitoolsUnderstandSoftware, ), we perform static analysis to identify Python classes. Class extraction is carried out using Python’s ast library, parsing all .py files to extract class skeletons while preserving class and method signatures and docstrings. Post-processing involves removing problematic classes (e.g., those with Python 2 syntax), ensuring only parsable and valid skeletons remain. From these, we randomly sample 200 classes per version. An example is shown in Table 1.

### v2

@@ Abstract.
Large language models (LLMs) have demonstrated strong performance on function-level code generation benchmarks, yet real-world software development increasingly demands class-level implementations that integrate multiple methods, attributes, and dependencies within authentic project contexts. This gap between benchmark performance and practical utility raises critical questions about LLMs’ readiness for production code assistance, particularly regarding their ability to generalize across familiar and novel codebases.
We introduce a benchmark derived from real-world open-source repositories, comprising classes divided into seen and unseen partitions to evaluate generalization under practical conditions. We systematically examine how input specification completeness and retrieval-augmented generation affect class-level correctness across multiple state-of-the-art LLMs.
Our evaluation reveals a substantial performance gap: while LLMs achieve 84–89% correctness on synthetic benchmarks, they attain only 25–34% on real-world class tasks, with minimal distinction between familiar and novel codebases. Comprehensive documentation provides marginal improvements (1–3%), whereas retrieval augmentation yields greater gains (4–7%) by supplying concrete implementation patterns. Error analysis identifies AttributeError, TypeError, and AssertionError as dominant failure modes, with distinct patterns between synthetic and real-world scenarios.
These findings provide actionable insights for enhancing context modelling, documentation strategies, and retrieval integration in production code assistance tools.
@@ 1. Introduction
Large language models (LLMs) have demonstrated remarkable progress in code generation, enabling tools like GitHub Copilot (githubGitHubCopilot) and Amazon CodeWhisperer (amazonCodeWhispererBecoming) to assist developers in writing their code on various levels of granularity (chen2021evaluating; tao2025retrieval). However, existing benchmarks such as CoderEval (yu2024codereval) and ClassEval (du2023classeval) primarily focus on either function-level generation or manually crafted class-level tasks. These benchmarks do not fully capture the complexities of real-world software projects, where developers frequently encounter interdependent classes, project-specific patterns, and the need to generalize to unseen codebases. The reliance on synthetic benchmarks creates a critical evaluation gap: if LLMs perform well on carefully constructed, self-contained tasks but fail on authentic production code with realistic dependencies and complexity, practitioners cannot reliably predict tool performance in actual development workflows. This disconnect between benchmark performance and real-world capability means developers may adopt LLM-based tools with inflated confidence, leading to productivity losses, increased debugging time, and potential integration of incorrect code into production systems. This gap raises a critical question: How well do LLMs generate class-level code for real-world projects, particularly when faced with unseen contexts? Addressing this question is essential for advancing LLM-assisted programming tools because it enables more accurate assessment of their practical capabilities, guides developers in setting appropriate expectations for tool reliability, and identifies specific technical limitations that must be overcome before these tools can be safely deployed in production environments where code correctness directly impacts software quality and development velocity.
@@
Current research on LLM-based code generation has largely centred on function-level tasks, with benchmarks like HumanEval (chen2021evaluating) and MBPP (austin2021program) evaluating standalone functions. While ClassEval (du2023classeval) introduced class-level evaluation, its manually crafted tasks do not reflect the intricacies of real-world projects. Specifically, synthetic benchmarks present several fundamental limitations: they typically feature self-contained classes with minimal external dependencies, whereas real-world code heavily relies on framework integrations, third-party libraries, and project-specific utilities; they use simplified test suites with straightforward assertions, while production code requires handling complex type systems, inheritance hierarchies, and dynamic attribute resolution; and they lack the evolutionary complexity of real codebases where classes must integrate with existing architectural patterns and legacy code. These differences mean that high performance on synthetic benchmarks does not guarantee success on realistic tasks, creating a false sense of LLM capabilities. Recent work on retrieval-augmented generation (RAG) for code (yang2025empirical; tao2025retrieval) has shown promise in improving LLM performance by incorporating relevant context, but its application to class-level generation in real-world settings remains underexplored. Furthermore, studies on repository-level code generation (tao2025retrieval) highlight the need for benchmarks that capture project-level dependencies, which are absent in existing evaluations.
@@
We address the issues above by curating a new benchmark of real-world class-level code from open-source projects, categorized as ‘seen’ (classes from repositories created before model training cutoffs, likely encountered during training) and ‘unseen’ (classes from repositories created after training cutoffs, guaranteeing they were not in training data). Our evaluation framework assesses LLM performance across different docstring variants, generation strategies, and RAG-augmented setups. We systematically analyze the impact of docstrings, the effectiveness of RAG, and the types of errors that arise in generated code, providing insights into the strengths and limitations of current LLMs. More specifically, we investigate the following research questions:
@@
  • RQ4 (Error Analysis): What are the most common errors in LLM-generated class-level code? We find that AttributeError, TypeError, and AssertionError account for 84% of all failures. Synthetic benchmarks show assertion-dominated failures while real-world code shows attribute-and-type-dominated failures, and RAG operates through error substitution, reducing logic failures while occasionally introducing dependency errors.
@@ 2.1. Large Language Models for Code Generation
Large Language Models (LLMs) have achieved remarkable success in code generation, enabling tools like GitHub Copilot and Amazon CodeWhisperer to assist developers in writing functions, classes, and even entire modules (chen2021evaluating; austin2021program). Recent works highlight the rapid evolution of LLMs for code-related tasks, including code completion, translation, and repair (jiang2024survey; bouzenia2024repairagent). Models such as GPT-4, CodeGen, and WizardCoder have demonstrated strong performance on benchmarks like HumanEval and MBPP, but their effectiveness in real-world scenarios—particularly for class-level code generation—remains underexplored (du2023classeval).
To understand the limitations of current LLMs, Wang et al. (wang2025towards) conducted an in-depth analysis of 557 incorrect code snippets generated by six prominent LLMs (CodeGen-16B, InCoder-1.3B, GPT-3.5, GPT-4, SantaCoder, and StarCoder) on the HumanEval dataset. They found that syntactic errors—such as missing or incorrectly structured code blocks—accounted for 40% of all errors, indicating that even advanced LLMs struggle with fundamental code structure. Furthermore, they identified common semantic errors, including incorrect logical flow and flawed conditional statements. Notably, GPT-4 exhibited only 7 of the 13 error categories, while smaller models displayed all or most error types, demonstrating a clear correlation between model size and error frequency. This analysis reveals that while LLMs have advanced significantly, they still face challenges in generating syntactically and semantically correct code, particularly for complex tasks. Complementing this, Yeo et al. (yeo2024framework) developed a systematic evaluation framework using LeetCode problems and found that GPT-4, when using optimal prompting strategies, outperforms 85% of human participants on coding contests. Their study revealed that prompt engineering significantly impacts LLM performance, with detailed prompts (including function signatures, input/output examples, and constraints) yielding substantially better results than basic descriptions. These findings underscore both the potential and limitations of LLMs: while they can match or exceed human performance on well-defined tasks with appropriate prompting, they struggle with structural correctness and semantic reasoning in more complex scenarios. Our study extends this understanding by evaluating LLMs on real-world class-level code generation, which presents additional challenges beyond the function-level tasks examined in prior work.
@@ 2.2. Benchmarks for Code Generation
Benchmarks serve as standardized evaluation datasets that measure LLM performance on code generation tasks, providing a crucial foundation for comparing model capabilities and tracking progress in the field. Existing benchmarks for code generation primarily focus on function-level tasks. For example, HumanEval (chen2021evaluating) and MBPP (austin2021program) evaluate standalone functions, while CoderEval (yu2024codereval) introduces non-standalone functions but still operates at the function level. ClassEval (du2023classeval) is the first benchmark to evaluate class-level code generation, but its tasks are manually crafted and do not reflect the complexities of real-world projects, such as cross-class dependencies and project-specific patterns.
Despite their widespread adoption, existing benchmarks face significant limitations in test coverage and real-world applicability. Liu et al. (liu2023your) introduced EvalPlus, which augments HumanEval with 80x more test cases using both LLM-based and mutation-based test generation strategies. Their extensive evaluation across 26 LLMs revealed that HumanEval’s original test suites suffer from severe test insufficiency, allowing models to pass with fragile code that fails on edge cases. Specifically, EvalPlus reduced pass@k scores by 19.3–28.9%, exposing previously undetected errors and even altering model rankings—for instance, WizardCoder-CodeLlama and Phind-CodeLlama outperformed ChatGPT on HumanEval+ but not on the original HumanEval. This finding demonstrates that prior benchmarks significantly overestimate LLM capabilities due to weak test suites. Building on this, Yu et al. (yu2024humaneval) proposed HumanEval Pro and MBPP Pro, which evaluate LLMs on self-invoking code generation—tasks requiring models to generate and subsequently utilize their own generated functions to solve more complex problems. Their evaluation of 20 LLMs revealed substantial performance degradation: for example, o1-mini achieved 96.2% pass@1 on HumanEval but only 76.2% on HumanEval Pro, indicating that while frontier LLMs excel at generating isolated code snippets, they struggle with reasoning about and reusing their own code. These studies collectively highlight critical gaps in current benchmarking practices: insufficient test coverage, limited evaluation of code reusability, and a focus on synthetic rather than real-world tasks. Our benchmark addresses these gaps by curating tasks from real-world repositories with comprehensive test suites, cross-class dependencies, and project-specific patterns that better reflect practical software development challenges.
@@ 2.3. Retrieval-Augmented Generation for Code
Retrieval-Augmented Generation (RAG) has emerged as a promising approach to enhance LLM performance by incorporating relevant context during generation (yang2025empirical; tao2025retrieval). Recent studies have explored RAG for code generation, demonstrating its potential to improve accuracy and contextual relevance (li2025coderag). For instance, CodeRAG (li2025coderag) integrates retrieval into the reasoning process, allowing LLMs to interact with source code snippets from repositories. However, most RAG studies focus on function-level or file-level tasks, and their application to class-level generation in real-world settings remains limited (tao2025retrieval).
To systematically investigate when and how retrieval benefits code generation, Wang et al. (wang2024coderag) introduced CodeRAG-Bench, a comprehensive benchmark encompassing three categories of tasks: basic programming (e.g., HumanEval), open-domain programming (e.g., ODEX), and repository-level problems. They aggregated retrieval sources from five diverse datastores—competition solutions, online tutorials, library documentation, StackOverflow posts, and GitHub repositories—to evaluate 10 retrievers and 10 LLMs. Their key findings reveal that while high-quality retrieved contexts can significantly improve code generation (e.g., GPT-4 gains substantial improvements on DS-1000 and repository-level tasks), current retrieval models struggle to fetch useful contexts, especially when queries have limited lexical overlap with relevant documents. Moreover, generation models face challenges in effectively utilizing retrieved contexts due to limited context window sizes and insufficient RAG integration capabilities. This analysis demonstrates that RAG’s potential for code generation remains underutilized, particularly for tasks requiring deep contextual understanding. Complementing this, Li et al. (li2025coderag) proposed a repository-level RAG framework called CodeRAG, which constructs dual graphs—a requirement graph modeling functional dependencies and a DS-code graph capturing both dependency and semantic relationships—to comprehensively retrieve supportive code snippets. Their agentic approach enables LLMs to iteratively reason and retrieve relevant code, achieving substantial improvements (40.90 and 37.79 Pass@1 increases for GPT-4o and Gemini-Pro on DevEval) compared to non-RAG baselines, and even outperforming commercial tools like GitHub Copilot and Cursor. These studies collectively demonstrate that RAG significantly enhances LLM performance on complex, repository-level tasks by providing contextually relevant information. However, the application of RAG to real-world class-level code generation, particularly in distinguishing between seen (classes from training projects) and unseen (classes from new projects) contexts, remains an open research question. Our study addresses this gap by evaluating RAG’s effectiveness on real-world class-level tasks and investigating whether retrieval strategies differ in their utility across seen versus unseen project contexts.
@@ 2.4. Docstrings and Prompt Engineering
Docstrings play a crucial role in code generation, as they provide natural language descriptions of code functionality. Studies have shown that the quality and presence of docstrings significantly impact LLM performance (yang2024less; ronanki2025prompt). For example, ClassEval (du2023classeval) includes human-labelled docstrings to mitigate data leakage and improve prompt quality. Recent work on docstring compression (yang2024less) further highlights the importance of concise and informative docstrings for effective code generation. Prompt engineering techniques, such as chain-of-thought and few-shot prompting (yuenprompting), have also been explored to enhance LLM performance in code-related tasks (tony2025prompting; sahoo2024systematic).
Investigating the role of docstring quality in code generation, Yang et al. (yang2024less) proposed ShortenDoc, a novel compression method specifically designed for docstrings in code generation tasks. Recognizing that docstrings often contain redundant information that increases token costs and processing time, they conducted extensive experiments on six datasets using five open-source LLMs (1B–10B parameters) and GPT-4o. Their findings demonstrate that state-of-the-art generic prompt compression methods achieve only  10% reduction before causing performance degradation, whereas ShortenDoc achieves 25–40% compression while preserving code generation quality. This result indicates that redundancy in docstrings can be strategically reduced without sacrificing LLM understanding, thereby improving efficiency and reducing API costs—particularly important for large-scale deployments. The study reveals that not all information in docstrings contributes equally to code generation quality, and careful compression can eliminate noise while retaining essential semantic content. Expanding on the broader role of prompt engineering, Taherkhani et al. (taherkhani2024epic) introduced EPiC (Evolutionary Prompt engineering for Code), which leverages a lightweight evolutionary algorithm to automatically refine prompts for code generation. Unlike iterative feedback methods that require numerous LLM interactions, EPiC achieves up to 6% improvement in pass@k while being 2–10 times more cost-effective than baselines. Their approach systematically optimizes prompt phrasing, structure, and specificity using minimal LLM queries, demonstrating that prompt engineering can substantially enhance code generation without extensive computational overhead. These findings collectively emphasize that both the content (what information docstrings contain) and presentation (how prompts are structured) critically influence LLM performance. While existing work has explored docstring quality and prompt optimization for function-level tasks, their impact on real-world class-level code generation—where docstrings must describe complex interdependencies and longer code structures—remains underexplored. Our study investigates how different docstring qualities (skeleton-only vs. full docstrings) affect LLM performance on class-level tasks, and whether this impact varies between seen and unseen project contexts.
@@ 3.1. Dataset Curation
To evaluate the performance of LLMs on real-world class-level code generation, we introduce RealClassEval, a dataset comprising two versions of Python classes: pre-cutoff and post-cutoff. Each version contains 200 classes mined from real-world GitHub open-source projects. The complete data curation process is illustrated in Figure 1, which shows our systematic approach from initial data collection through final dataset construction.
3.1.1. Data Collection:
@@
The pre-cutoff data is sourced from the CodeSearchNet dataset (husain2019codesearchnet), which is one of the most widely used datasets for training code-related LLMs. CodeSearchNet has been used as a pre-training corpus for numerous prominent models including StarCoder (li2023starcoder), CodeT5+ (wang2023codet5+), and UniXcoder (guo2022unixcoder), and has been used in numerous existing research papers (jiang2024survey; shi2023cocosoda; rahman2024automatic; orel2025codet). The dataset comprises 2 million (comment, code) pairs from open-source GitHub repositories across six programming languages. Even though we cannot guarantee that this dataset was indeed part of the training data of the LLMs we investigated in this research, we believe that it is a safe assumption to make, given the widespread adoption and popularity of this dataset in the LLM training community. The post-cutoff data consists of repositories created after March 31, 2025  [^1] All LLMs under investigation were either released before this date or their knowledge cut-off dates are before this date., guaranteeing these classes are unseen by the evaluated LLMs.
For the post-cutoff data collection, we used the GitHub search API to systematically identify Python repositories created after March 31, 2025. We selected repositories that: (1) were created after our cutoff date to ensure temporal separation from LLM training data, (2) had active development with at least 10 commits to ensure maturity, (3) contained Python as the primary programming language, and (4) were not forks to eliminate duplicates. This systematic selection process yielded a pool of candidate repositories from which we subsequently applied filtering criteria to identify engineered projects.
3.1.2. Data Filtering:
To ensure our dataset comprises high-quality, production-level software, we retain only engineered projects (munaiah2017curating) from both the CodeSearchNet dataset and the post-cutoff repositories, adhering to the filtering criteria outlined in (xiao2025self). Importantly, we did not include all projects from CodeSearchNet; rather, we applied the same rigorous filtering criteria to both data sources to ensure consistency and quality. Specifically, we applied three filtering criteria to both pre-cutoff (CodeSearchNet) and post-cutoff (newly collected) repositories:
  1. (1)
    License Declaration: We excluded repositories not declaring a license or using non-standard licenses (marked as “Other” in the GitHub search tool). For the remaining repositories, we removed projects declaring licenses not commonly used for software projects, such as Creative Commons Attribution licenses and font licenses, as these typically indicate documentation or non-code projects.
  2. (2)
    Project Size and Activity: We filtered out repositories with fewer than a minimum threshold of stars (10+) and commits (50+) to ensure the projects represent mature, community-recognized software rather than personal experiments or abandoned prototypes.
  3. (3)
    Code Quality Indicators: We excluded projects with incomplete documentation, lack of structured code organization (e.g., no clear module structure), or repositories that primarily contain tutorials, educational materials, or code snippets rather than production-quality software.
These filtering criteria, applied uniformly to both CodeSearchNet and post-cutoff repositories, ensure that our dataset comprises classes from well-maintained, professionally developed open-source projects that reflect real-world software engineering practices.
3.1.3. Static Analysis and Class Extraction:
From the filtered engineered projects in both data sources, we use Understand™ (scitoolsUnderstandSoftware) to perform static analysis and identify Python classes. Class extraction is carried out using Python’s ast library, parsing all .py files to extract class skeletons while preserving class and method signatures and docstrings. An example of the class skeleton extraction process is shown in Table 1, which illustrates how we transform a complete human-written class into a skeleton that serves as input for LLM code generation.
3.1.4. Post-processing and Sampling:
Post-processing involves removing problematic classes (e.g., those with Python 2 syntax), ensuring only parsable and valid skeletons remain. Following ClassEval’s exclusion criteria (du2023classeval), we omit tasks with complex execution dependencies, identified through static metrics. However, unlike ClassEval, we do not filter by domain, retaining a diverse range of topics such as Game Development, File Handling, and Management Systems. This approach ensures our dataset reflects the heterogeneity of real-world projects, providing a robust foundation for evaluating LLM performance on class-level code generation. From the remaining classes, we randomly sample 200 classes per version (pre-cutoff and post-cutoff) to create a balanced dataset for evaluation.